# Personalizing Student-Agent Interactions Using Log-Contextualized Retrieval Augmented Generation (RAG)

### v1

@@ Abstract
ning and critical thinking. This is essential for adapting pedagogical agents to students’ learning and problem-solving skills in STEM+C settings. While large language models (LLMs) facilitate dynamic pedagogical interactions, potential hallucinations can undermine confidence, trust, and instructional value. Retrieval-augmented generation (RAG) grounds LLM outputs in curated knowledge, but its effectiveness depends on clear semantic links between user input and a knowledge base, which are often weak in student dialogue. We propose log-contextualized RAG (LC-RAG), which enhances RAG retrieval by incorporating environment logs to contextualize collaborative discourse. Our findings show that LC-RAG improves retrieval over a discourse-only baseline and allows our collaborative peer agent, Copa, to deliver relevant, personalized guidance that supports students’ critical thinking and epistemic decision-making in a collaborative computational modeling environment, C2STEM.
@@ 1 Introduction
The integration of generative AI (GenAI) into collaborative learning environments raises important questions about how this technology can best help students construct knowledge, make decisions, and engage in shared problem-solving activities. Collaborative computational modeling — where students use code to construct models that simulate and analyze scientific processes [34] — enhances science, technology, engineering, mathematics, and computing (STEM+C) learning [19]. However, these complex environments can also create problem-solving challenges [2, 12], highlighting the need for adaptive supports that align with students’ evolving understanding and promote epistemic agency.
Vygotsky’s zone of proximal development (ZPD) [39] emphasizes guided support as paramount in helping learners accomplish tasks they cannot yet complete independently [8]. Pedagogical agents help bridge this gap by guiding individual learners and facilitating collaboration [8, 37, 33]. However, developing agents to support collaborative STEM+C problem-solving is non-trivial, as agents must adapt to both individual learners and groups while guiding learning and collaboration across multiple domains simultaneously (e.g., physics and computing).
Collaborative dialogue provides insights into students’ learning and critical thinking, helping identify knowledge gaps and problem-solving approaches [34]. This information can improve pedagogical agents, which struggle to (1) capture the nuances of individual contributions in collaborative settings; (2) ground interactions in ongoing problem-solving discussions; and (3) offer dynamic and personalized adaptive scaffolding beyond rigid, rule-based approaches.
However, dialogue alone is insufficient for effectively guiding students through challenges. Many students struggle with help-seeking and often fail to recognize when they need assistance [1, 32]. Jurenka et al. (2024) [20] emphasize that pedagogical agents should “see what the student sees,” highlighting the need for agents that are aware of students’ actions inside the learning environment. Previous research has shown that integrating collaborative discourse with learning environment activity logs can provide a more comprehensive understanding of student learning [12, 35, 36]. In this context, large language models (LLMs) offer a promising solution [37].
While LLMs have been applied across various artificial intelligence in education (AIED) tasks like intelligent tutoring [15, 37] and formative assessment scoring [6, 5, 10], they are prone to hallucinations that lead to errors and undermine trust in educational systems [18, 9]. Fine-tuning can help reduce hallucinations [21], but it typically requires substantial amounts of data and computing power [9] and may limit LLMs’ ability to generalize beyond their training [14].
@@ 2 Background
RAG is valuable for aligning LLM outputs with educator preferences, giving teachers greater control over responses. It has been effective for tasks like intelligent tutoring [23], data visualization [41], MOOC recommendations [31], question answering [27], and short-answer scoring [13]. However, RAG requires a semantic link between input text and the knowledge base, which is often missing in collaborative discourse. Research suggests that integrating log and discourse data via LLM-based summarization could help address this gap [12, 11].
While several researchers have identified this disconnect, few have explored solutions. The MemoRAG framework [30] utilizes a lightweight LLM as a memory mechanism to create a compressed representation of the knowledge base, generating ‘clues’ from user input to improve knowledge base alignment. However, training such a memory module can be impractical in educational settings due to high computational demands and limited data [4].
@@
Tools like LlamaIndex’s Query Transformations module [25] enable users to refine queries by generating a hypothetical document or answer for RAG retrieval. However, this assumes the initial query is semantically relevant to the task and knowledge base, which is often not the case in collaborative discourse due to knowledge gaps and off-topic conversation. A notable research gap exists in developing RAG methods that effectively retrieve relevant knowledge during collaborative dialogue, which we address through log-contextualized summarization using LC-RAG.
@@ 3.1 LC-RAG
LC-RAG augments RAG retrieval by integrating student interactions and log data for use in pedagogical agents. Instead of using raw text for semantic search, LC-RAG utilizes an LLM to generate summaries that combine discourse and log data, contextualizing student interactions. These summary embeddings are mapped to the knowledge base, enabling contextualized domain knowledge retrieval during LLM inference. Figure 1 illustrates this process. LC-RAG supports students’ epistemic agency by grounding agent responses in contextually relevant knowledge, helping them construct, validate, and apply domain concepts during collaborative problem-solving. The following subsections discuss the design and implementation of LC-RAG in this paper.
@@
3.2 C2STEM, Study Participants, and Data
C2STEM [19] is a collaborative computational modeling environment targeting physics and computing where students create and debug computational models simulating scientific phenomena (see Figure 2). In this study, students engaged with a one-dimensional Truck Task activity where they modeled a truck’s motion accelerating from rest, cruising, and decelerating to stop at a stop sign.
Figure 2: Example solution for the C2STEM Truck Task with task context categories.
To answer RQ1, we conducted an eight-week study with 24 high school sophomores, ages 15-16, participating in a kinematics curriculum with one two-hour session each week. Students worked in pairs, sharing a laptop. We collected data on 18 consented students during the Truck Task, resulting in 2,275 logged actions and 2,786 spoken utterances over nine hours of collaborative discourse. Audio recordings were transcribed using Otter.ai, corrected by two research team members and aligned with timestamped actions in the C2STEM environment.
We segmented the multimodal data based on environment logs related to the model components students worked on. Using an abstract syntax tree (AST [16]), we classified logged student actions into four task context categories [36]: (1) Initializing Variables, (2) Updating Variables, Each Simulation Step, (3) Conditional Statements, and (4) Updating Variables, Under Conditions. Figure 2 shows these categories in a sample student model. For instance, blocks placed under “When Green Flag Clicked” were classified as Initializing Variables segments. This method ensures that student discourse reflects their problem-solving context. We identified a total of n=216 problem-solving segments, with Conditional Statements being the most common (77 segments, average discourse length of 790 characters) and Updating Variables, Each Simulation Step the least common (36 segments, 443 characters).
@@ 3.2 C2STEM, Study Participants, and Data
To answer RQ2, we conducted two studies with a total of n=18 students interacting with Copa: (1) a one-day, 2.5-hour classroom study with 16 high school freshmen (ages 14–15), where students solved a simplified version of the Truck Task (excluding deceleration and stopping conditions); and (2) two one-hour focus groups with six students — four from the classroom study and two from the earlier study for RQ1 — who debugged a fictional student’s faulty Truck Task model using Copa.
Classroom study participants took part in the research as part of their regular curriculum. The study began with an instructional session introducing students to C2STEM, followed by the Truck Task problem-solving activity using Copa. After the activity, one of this paper’s authors facilitated an interactive class discussion using a Miro [^2] https://miro.com/ storyboarding exercise to reflect on Copa’s accuracy, helpfulness, and trustworthiness. The session concluded with a semi-structured discussion, led by another author, exploring students’ broader perceptions of Copa. Data collected included student conversations, environment actions, agent interactions, screen recordings, Miro board notes, and discussion notes.
Focus group participants volunteered to participate outside of regular class time and received a \$25 Amazon gift card for their involvement. Students were first given task instructions explaining that they would use Copa to debug a fictional student’s erroneous C2STEM Truck Task code. After completing the debugging activity, one of this paper’s authors led a semi-structured discussion to gather students’ perceptions of Copa. Participants also completed a survey about their experience, including their likes and dislikes, as well as their views on the accuracy, helpfulness, and trustworthiness of the agent’s feedback. Data collected included student conversations, webcam recordings, environment actions, agent interactions, screen recordings, discussion notes, and survey responses.
All studies were conducted in the southeastern United States with approval from Vanderbilt University’s Institutional Review Board (IRB), and all participating students and their parents provided assent and consent, respectively. Demographic data was only able to be collected for the focus groups: 4 freshmen (no prior C2STEM experience), 2 seniors (with C2STEM experience); 4 female, 1 male, 1 self-unidentified; 2 Hispanic, 2 White, 1 Black, 1 Persian.
3.3 Experimental Design
We evaluate LC-RAG’s ability to retrieve relevant information from a knowledge base relative to a discourse-only baseline (i.e., without using environment logs). Our analysis involves five embedding models of different dimensions (discussed shortly) to assess LC-RAG’s performance in terms of relevance, accuracy, and helpfulness (RQ1) across 216 student problem-solving segments. We also examine performance across task context categories and conduct classroom and focus group studies with 18 students problem-solving in C2STEM using Copa (RQ2). For our knowledge base, we selected two high school textbooks — one for each domain in C2STEM, physics [38] and computing [3] — due to their open-source availability, curriculum relevance, and age appropriateness.
@@ 3.3 Experimental Design
While previous AIED research has effectively used LLM’s and LLM-as-a-Judge for evaluation [29, 27], studies show that variations in prompt phrasing [28] and example ordering [26] can significantly impact outcomes. To enhance robustness, Mizrahi et al. (2024) [28] recommend multi-prompt evaluation, aggregating results across multiple templates. Accordingly, we created three separate evaluation prompts, aggregating the judge’s predictions based on majority vote. Chunk order was randomized each run pursuant to Lu et al. (2021) [26]. All three evaluation prompts are available in our Supplementary Materials.
@@
To validate our LLM-as-a-Judge protocol, we randomly sampled 80 segments across all model-dimension combinations. Two authors scored all 80 retrieval pairs via consensus coding [42] and compared LC-RAG’s retrievals to the baseline using the human-crafted prompt. We assessed inter-rater reliability (IRR) using Cohen’s Quadratic Weighted Kappa (QWK) [7] due to the data being ordinal, resulting in QWK$=0.49$ (moderate agreement). Qualitatively, both the LLM and humans prioritized the inclusion of relevant physics and computing concepts related to C2STEM in their justifications. Some disagreements were expected due to the subjectivity of relevance and helpfulness evaluation.
@@
The four task context categories discussed in Section 3.2 require different domain knowledge. For example, Initializing Variables primarily involves initializing constants and variables (computing), while Updating Variables, Each Simulation Step requires an understanding of loops (computing) and kinematic variable relationships (physics). This variation necessitated evaluating LC-RAG’s performance across these contexts (RQ1), which we addressed by averaging LC-RAG and baseline win rates across the five embedding models for all four task context categories to identify trends. For RQ1, we sampled LLM responses from all three prompt templates for qualitative analysis, enhancing our quantitative findings with deeper insights.
@@
To determine Copa’s epistemic value to students and its ability to promote critical thinking (RQ2), we conducted a qualitative analysis by reviewing students’ interactions with Copa (including retrieved domain knowledge), Miro boards, discussion notes, and focus group survey responses, memoing key findings across all sources [17] and aggregating the results. Results for all three RQs are presented in the next section.
@@ 4.1 LC-RAG (RQ1)
RQ1 asked, “How does LC-RAG compare to a discourse-only baseline in terms of the relevance, accuracy, and helpfulness of retrieved knowledge, and how does its performance vary across task context categories?,” which we evaluated using our LLM-as-a-Judge protocol across five embedding model-dimension combinations (see Section 3.3). Results are presented in Table 1.
@@
LC-RAG outperformed the baseline in 4 out of 5 embedding model-dimension combinations. Qualitatively, both approaches retrieved information relevant to kinematics and computing. However, LC-RAG achieved more precise retrieval, isolating a smaller set of relevant chunks compared to the baseline. LC-RAG retrieved 167 unique chunks from a knowledge base of 7,386, while the baseline retrieved 267. This advantage was especially clear when student discussions were off-topic. For example, consider a segment where students said, “S1: We need help. S2: No we don’t. We got this. We can do this.” In this case, using the voyage-3-lite embedding model, the baseline retrieved an irrelevant chunk about a ‘90s TV show (“…from an episode of Ren and Stimpy…”), while LC-RAG retrieved a chunk related to scientific modeling.
The baseline outperformed LC-RAG with OpenAI’s “small” embedding model due to the general nature of the knowledge base, which comprised physics and computing domain knowledge outside the scope of the C2STEM Truck Task (e.g., thermodynamics and memory management). The baseline’s raw discourse typically lacked references to the C2STEM environment or Truck Task, whereas LC-RAG’s integrated discourse and log summaries (see Section 3.1) reframed student conversations within the context of C2STEM, the Truck Task, and their environment actions. While this contextualization made LC-RAG summaries more meaningful within C2STEM, it hindered retrieval performance, as the knowledge base lacked C2STEM or Truck Task information. Although the baseline outperformed LC-RAG quantitatively with this model, qualitative analysis revealed that both were prone to irrelevant retrievals, suggesting the limitation stems from the untargeted knowledge base rather than the retrieval mechanism itself. We address this while answering RQ2.
@@
Across all four task context categories (and overall), LC-RAG achieved a higher win rate than the baseline when considering results from all five embedding models. LC-RAG’s outperformance was more pronounced in Initializing Variables (45% for LC-RAG, 33% for the baseline) and Updating Variables, Each Simulation Step (46% for LC-RAG, 38% for the baseline) segments but less pronounced for Conditional Statements and Updating Variables, Under Conditions, differing only by 1-2 percentage points.
Performance gains were reduced for Updating Variables, Under Conditions and Conditional Statements for the same reason as previously discussed: when using the OpenAI “small” model, the LC-RAG summaries were too task-specific to align well with the generalized knowledge base. Excluding this model, all four task-context categories showed clear improvement over the baseline. This underperformance, while limited to a single embedding model, highlights a limitation in how our log-derived task context categories contextualize student discourse. Our RQ2 analysis explores an alternative approach to integrating environment logs that generates segment summaries with stronger semantic links to the knowledge base.

### v2

@@ Abstract
ning and critical thinking, which is essential for personalizing pedagogical agent interactions in STEM+C settings. While large language models (LLMs) facilitate dynamic pedagogical interactions, hallucinations undermine confidence, trust, and instructional value. Retrieval-augmented generation (RAG) grounds LLM outputs in curated knowledge but requires a clear semantic link between user input and a knowledge base, which is often weak in student dialogue. We propose log-contextualized RAG (LC-RAG), which enhances RAG retrieval by using environment logs to contextualize collaborative discourse. Our findings show that LC-RAG improves retrieval over a discourse-only baseline and allows our collaborative peer agent, Copa, to deliver relevant, personalized guidance that supports students’ critical thinking and epistemic decision-making in a collaborative computational modeling environment, C2STEM.
@@ 1 Introduction
The integration of generative AI (GenAI) into collaborative learning environments raises important questions about how this technology can best help students construct knowledge, make decisions, and engage in shared problem-solving activities. Collaborative computational modeling — where students use code to construct models that simulate and analyze scientific processes [22] — enhances science, technology, engineering, mathematics, and computing (STEM+C) learning [12]. However, these complex environments can also create problem-solving challenges [9], highlighting the need for adaptive supports that align with students’ evolving understanding and promote epistemic agency.
Vygotsky’s zone of proximal development (ZPD) [27] emphasizes guided support as paramount in helping learners accomplish tasks they cannot yet complete independently [5]. Pedagogical agents help bridge this gap by guiding individual learners and facilitating collaboration [5, 25]. However, developing agents to support collaborative STEM+C problem-solving is non-trivial, as agents must adapt to both individual learners and groups while guiding learning and collaboration across multiple domains simultaneously (e.g., physics and computing).
Collaborative dialogue provides insights into students’ learning and critical thinking, helping identify knowledge gaps and problem-solving approaches [22]. This information can improve pedagogical agents, which struggle to (1) ground their interactions in ongoing problem-solving discussions; and (2) offer dynamic and personalized adaptive scaffolding beyond rigid, rule-based approaches. However, dialogue alone is insufficient for effectively guiding students through challenges, as many students struggle with help-seeking and often fail to recognize when they need assistance [21].
Jurenka et al. (2024) [13] emphasize that pedagogical agents should “see what the student sees,” highlighting the need for agents that are aware of students’ actions inside the learning environment. Previous research has shown that integrating collaborative discourse with learning environment activity logs can provide a more comprehensive understanding of student learning [9, 24, 23]. In this context, large language models (LLMs) offer a promising solution [25].
While LLMs have been applied across various artificial intelligence in education (AIED) tasks like intelligent tutoring [25] and formative assessment scoring [4, 3, 7], they are prone to hallucinations that lead to errors and undermine trust in educational systems [6]. Fine-tuning can help reduce hallucinations, but it typically requires substantial amounts of data and computing power and may limit LLMs’ ability to generalize beyond their training [10].
@@ 2 Background
RAG is valuable for aligning LLM outputs with educator preferences, giving teachers greater control over responses. However, it requires a semantic link between input text and the knowledge base, which is often missing in collaborative discourse. While several researchers have identified this disconnect, few have explored solutions.
The MemoRAG framework [20] utilizes a lightweight LLM as a memory mechanism to create a compressed representation of the knowledge base, generating ‘clues’ from user input to improve knowledge base alignment. However, training such a memory module can be impractical in educational settings due to high computational demands and limited data [2].
@@
LlamaIndex’s Query Transformations module [16] enables users to refine queries by generating a hypothetical document or answer for RAG retrieval. However, this assumes the initial query is semantically relevant to the task and knowledge base, which is often not the case in collaborative discourse due to knowledge gaps and off-topic conversation. A notable research gap exists in developing RAG methods that effectively retrieve relevant knowledge during collaborative dialogue, which we address through log-contextualized summarization using LC-RAG.
@@ 3.1 LC-RAG
LC-RAG augments RAG retrieval by integrating student interactions and environment log data for use in pedagogical agents. Instead of using raw text for semantic search, LC-RAG utilizes an LLM to generate summaries that combine discourse and log data, contextualizing student interactions. These summary embeddings are mapped to the knowledge base, enabling contextualized domain knowledge retrieval during LLM inference. Figure 1 illustrates this process. LC-RAG supports students’ epistemic agency by grounding agent responses in contextually relevant information, helping them construct, validate, and apply disciplinary knowledge during collaborative problem-solving. Section 3.3 discusses the design and implementation of LC-RAG in this paper.
@@
3.2 C2STEM Learning Environment
C2STEM [12] is a collaborative computational modeling environment targeting physics and computing where students create and debug computational models simulating scientific phenomena (shown in Figure 2). In this study, students engaged with a one-dimensional Truck Task activity where they modeled a truck’s motion accelerating from rest, cruising, and decelerating to stop at a stop sign.
Figure 2: C2STEM Truck Task example solution with task context categories.
3.3 Experimental Design
All studies described below were conducted in Nashville, TN, USA with approval from Vanderbilt University’s Institutional Review Board (IRB), and all participating students and their parents provided assent and consent, respectively.
3.3.1 RQ1 (LC-RAG)
To answer RQ1, we conducted an eight-week study with 24 high school sophomores, ages 15-16, participating in a kinematics curriculum with one two-hour session each week. Students worked in pairs, sharing a laptop. We collected (and transcribed) audio and environment log data on 18 consented students during the Truck Task, resulting in 2,786 spoken utterances and 2,275 logged actions over nine hours of collaborative discourse.
We used environment logs to segment the multimodal data based on the model components students worked on. Using an abstract syntax tree (AST), we classified logged student actions into four task context categories [23]: (1) Initializing Variables, (2) Updating Variables, Each Simulation Step, (3) Conditional Statements, and (4) Updating Variables, Under Conditions (see Figure 2). For instance, blocks placed under “When Green Flag Clicked” were labeled as Initializing Variables segments. This method ensured that student discourse reflected their problem-solving context. We identified a total of n=216 problem-solving segments, with Conditional Statements being the most common (77 segments, average discourse length of 790 characters) and Updating Variables, Each Simulation Step the least common (36 segments, 443 characters).
@@ 3.3.1 RQ1 (LC-RAG)
We evaluated LC-RAG’s ability to retrieve relevant information from a knowledge base relative to a discourse-only baseline (i.e., without using environment logs). Our analysis involved five embedding models of different dimensions (discussed shortly) to assess LC-RAG’s performance in terms of relevance, accuracy, and helpfulness across the 216 student problem-solving segments. For our knowledge base, we selected two high school textbooks — one for each domain in C2STEM, physics [26] and computing [1] — due to their open-source availability, curricular relevance, and age appropriateness.
@@
While previous AIED research has effectively used LLM-as-a-Judge for evaluation [19], studies show that variations in prompt phrasing [18] and example ordering [17] can significantly impact outcomes. To enhance robustness, Mizrahi et al. (2024) [18] recommend multi-prompt evaluation, aggregating results across multiple templates. Accordingly, we created three separate evaluation prompts, aggregating the judge’s predictions based on majority vote. Chunk order was randomized each run pursuant to Lu et al. (2021) [17]. All three evaluation prompts are available in our Supplementary Materials.
@@
To validate our LLM-as-a-Judge protocol, we randomly sampled 80 segments across all model-dimension combinations. Two authors scored all 80 retrieval pairs via consensus coding and compared LC-RAG’s retrievals to the baseline’s using the same criteria given to the LLM. We assessed inter-rater reliability (IRR) using Cohen’s Quadratic Weighted Kappa (QWK) due to the data being ordinal, resulting in QWK$=0.49$ (moderate agreement).
@@
Each of the four task context categories discussed previously requires different domain knowledge. For example, Initializing Variables primarily involves initializing constants and variables (computing), while Updating Variables, Each Simulation Step requires an understanding of loops (computing) and kinematic variable relationships (physics). This variation necessitated evaluating LC-RAG’s performance across these contexts, which we addressed by averaging LC-RAG and baseline win rates across the five embedding models for all four task context categories to identify trends. For RQ1, we sampled LLM responses from all three prompt templates for qualitative analysis, enhancing our quantitative findings with deeper insights.
3.3.2 RQ2 (Copa)
To answer RQ2, we conducted two studies with a total of n=18 students interacting with Copa: (1) a one-day, 2.5-hour classroom study with 16 high school freshmen (ages 14–15), where students solved a simplified version of the Truck Task (excluding the deceleration and stopping conditions); and (2) two one-hour focus groups with six students — four from the classroom study and two from the earlier study for RQ1 — who debugged a fictional student’s faulty Truck Task model using Copa.
Classroom study participants took part in the research as part of their regular curriculum. The study began with an instructional session introducing students to C2STEM, followed by the Truck Task problem-solving activity using Copa. After the activity, one of this paper’s authors facilitated an interactive class discussion using a Miro storyboarding exercise to reflect on Copa’s accuracy, helpfulness, and trustworthiness. The session concluded with a semi-structured discussion, led by another author, exploring students’ broader perceptions of Copa. Data collected included student conversations, environment actions, agent interactions, screen recordings, Miro board notes, and discussion notes.
Focus group participants volunteered to participate outside of regular class time and received a \$25 Amazon gift card for their involvement. Students were first given task instructions explaining that they would use Copa to debug a fictional student’s erroneous C2STEM Truck Task code. After completing the debugging activity, one of this paper’s authors led a semi-structured discussion to gather students’ perceptions of Copa. Participants also completed a survey about their experience, including their likes and dislikes as well as their views on the accuracy, helpfulness, and trustworthiness of the agent’s feedback. Data collected included student conversations, webcam recordings, environment actions, agent interactions, screen recordings, discussion notes, and survey responses.
@@ 3.3.2 RQ2 (Copa)
To determine Copa’s epistemic value to students and its ability to promote critical thinking, we conducted a qualitative analysis by reviewing students’ interactions with Copa (including retrieved domain knowledge), Miro boards, discussion notes, and focus group survey responses, memoing key findings across all sources [11] and aggregating the results. Results for both RQs are presented in the next section.
@@ 4.1 LC-RAG (RQ1)
We answer RQ1 by comparing LC-RAG’s retrieval performance to the baseline’s across five embedding model-dimension combinations (see Section 3.3.1). Results are presented in Table 1.
@@
LC-RAG outperformed the baseline in 4 out of 5 embedding model-dimension combinations, achieving more precise retrieval by isolating a small-
er set of relevant chunks compared to the baseline. LC-RAG retrieved 167 unique chunks from a knowledge base of 7,386, while the baseline retrieved 267. This was advantageous, particularly when student discussions were off-topic and resulted in irrelevant retrieval for the baseline. For example, consider a segment where students said, “S1: We need help. S2: No we don’t. We got this. We can do this.” In this case, using the voyage-3-lite embedding model, the baseline retrieved an irrelevant chunk about a ‘90s TV show (“…from an episode of Ren and Stimpy…”), while LC-RAG retrieved a chunk related to scientific modeling.
The baseline outperformed LC-RAG with OpenAI’s “small” embedding model due to the general nature of the knowledge base, which comprised physics and computing domain knowledge outside the scope of the C2STEM Truck Task (e.g., thermodynamics and memory management). The baseline’s raw discourse often lacked references to the C2STEM environment or Truck Task, whereas LC-RAG’s integrated discourse and log summaries (see Section 3.1) reframed student conversations within the context of C2STEM, the Truck Task, and their environment actions. While this contextualization made LC-RAG summaries more meaningful within C2STEM, it hindered retrieval performance, as the knowledge base lacked C2STEM or Truck Task information. Although the baseline outperformed LC-RAG quantitatively with this model, qualitative analysis revealed that both were prone to irrelevant retrievals, suggesting the limitation stems from the untargeted knowledge base rather than the retrieval mechanism itself. We address this while answering RQ2.
@@
Across all four task context categories (and overall), LC-RAG achieved a higher win rate than the baseline when considering results from all five embedding models. LC-RAG’s outperformance was more pronounced in Initializing Variables (45% for LC-RAG, 33% for the baseline) and Updating Variables, Each Simulation Step (46% for LC-RAG, 38% for the baseline) segments but less pronounced for Conditional Statements and Updating Variables, Under Conditions, differing only by 1-2%.
Performance gains were reduced for Updating Variables, Under Conditions and Conditional Statements for the same reason as previously discussed: when using the OpenAI “small” model, the LC-RAG summaries were too task-specific to align well with the generalized knowledge base. Excluding this model, all four task-context categories showed clear improvement over the baseline. This underperformance, while limited to a single embedding model, highlights a limitation in how our task context categories contextualize student discourse. Our RQ2 analysis explores an alternative approach to integrating log data that generates segment summaries with stronger semantic links to the knowledge base.